# Predicting Energy Budgets in Droplet Dynamics: A Recurrent Neural Network Approach

### v1

@@
Diego A. de Aguiar 111diego.aguiar@unesp.br, ORCID: $0$-0001-5159-867X Departamento de Matemática e Computação, Faculdade de Ciências e Tecnologia, Universidade Estadual Paulista “Júlio de Mesquita Filho”, Presidente Prudente, Brazil Hugo L. França 222franca.hugo1@gmail.com, ORCID: $0$-0002-5361-7704 Van der Waals-Zeeman Institute, Institute of Physics, University of Amsterdam, Amsterdam, The Netherlands Cassio M. Oishi333cassio.oishi@unesp.br, ORCID: $0$-0002-5361-7704 Departamento de Matemática e Computação, Faculdade de Ciências e Tecnologia, Universidade Estadual Paulista “Júlio de Mesquita Filho”, Presidente Prudente, Brazil
@@
Neural networks in fluid mechanics offer an efficient approach for exploring complex flows, including multiphase and free surface flows. The recurrent neural network, particularly the Long Short-Term Memory (LSTM) model, proves attractive for learning mappings from transient inputs to dynamic outputs. This study applies LSTM to predict transient and static outputs for fluid flows under surface tension effects. Specifically, we explore two distinct droplet dynamic scenarios: droplets with diverse initial shapes impacting with solid surfaces, as well as the coalescence of two droplets following collision. Using only dimensionless numbers and geometric time series data from numerical simulations, LSTM predicts the energy budget. The marker-and-cell front-tracking methodology combined with a marker-and-cell finite-difference strategy is adopted for simulating the droplet dynamics. In addition, as the code must handle various drop shapes, we have also employed a method designed for extracting particles from interfaces in digital images. Using a recurrent neural network (RNN) architecture fed with time series data derived from geometrical parameters, as for example droplet diameter variation, our study shows the accuracy of our approach in predicting energy budgets, as for instance the kinetic, dissipation, and surface energy trends, across a range of Reynolds and Weber numbers in droplet dynamic problems. Finally, a two-phase sequential neural network using only geometric data, which is readily available in experimental settings, is employed to predict the energies and then use them to estimate static parameters, such as the Reynolds and Weber numbers. While our methodology has been primarily validated with simulation data, its adaptability to experimental datasets is a promising avenue for future exploration. Finally, we hope that our strategy can be useful for diverse applications, spanning from inkjet printing to combustion engines, where the prediction of energy budgets or dissipation energies is crucial. This analysis provides insights for developing machine learning modeling using numerical and/or experimental data from droplet dynamics.
keywords: Droplets |normal-||| Numerical solution |normal-||| Prediction |normal-||| Surface Tension |normal-||| LSTM |normal-||| Energy budget
@@ I.   Introduction
Droplet dynamics, a phenomenon of remarkable complexity and practical significance, extends its influence across various domains of science and engineering, ranging from raindrop splashes on leaves to the precise deposition of microdroplets in advanced manufacturing processes. Achieving a fundamental understanding of the dynamics governing droplet impact represents a significant scientific challenge, needing the collaborative engagement of experimental, analytical and numerical methodologies. This concerted effort is essential to unravel the complexities inherent in the complex phenomenon of droplet impact dynamics.
Advancements in these methodologies have unveiled a wealth of intricate physical mechanisms governing liquid-drop impacts on rigid surfaces. These insights not only contribute to the fundamental understanding within the realm of fluid dynamics but also create innovations across a diverse array of applications. The implications of droplet impact dynamics reverberate through fields such as aerospace, energy, materials science, biotechnology, forensic analysis, and studies related to food and cosmetics, among others.
In recent decades, the exploration of the droplet dynamics has yielded a rich body of literature, including excellent review articles that provide comprehensive insights into the multifaceted aspects of droplet dynamics. Notable contributions in this context have been made, as well detailed in [1, 2, 3, 4]. These articles clarify some important aspects as for instance: the dimensionless numbers and their influence on the maximum spreading of the droplet, different regimes considering pre- and post-impact and their consequences in terms of energy mechanisms, etc.
Concerning droplet dynamics and its applications, the complexities of the inkjet printing process stand out as a pertinent challenge. This process involves various fluid dynamics challenges that significantly impact the outcome of the printing process. A comprehensive overview of these challenges, particularly focusing on fundamental aspects, has been extensively discussed in the review paper by [5]. A fundamental question of inkjet printing challenge lies in the collision dynamics of droplets during their trajectory from ejection to eventual impact on the substrate. The understanding of this collision process is important, influencing the spatial distribution, size, and morphology of the printed droplets. As droplets traverse through the air, interactions can be observed, governed by fluid dynamics principles, as for instance coalescence, fragmentation, bouncing and other complex behaviors. After this, a different dynamic happens when the droplets impact on the substrate. Factors such as droplet velocity, angle of impact, and the properties of the drop and substrate all contribute to the final result. Moreover, the subsequent spreading of the droplets upon contact with the substrate adds another layer of complexity to the overall process. The dynamics of spreading govern how the ink interacts with the substrate surface, influencing factors as for instance the shape of the droplet. In essence, the challenges within inkjet printing underscore the profound interplay of fluid dynamics phenomena, encompassing droplet collisions, substrate interactions, and spreading dynamics. In this work, we have focused on two stages described in the review paper [5]: the impact of a droplet with different shapes in a substrate and the coalescence of droplets during the collision process.
Over the past decades, the simulation of droplet dynamics has been adopted for advancing our understanding of complex fluid phenomena [6]. These numerical investigations have opened new horizons, allowing for innovative approaches in the study of droplet behaviors and interactions. In these computational frameworks, the governing equations are rigorously solved on fixed or adaptive meshes, employing classical discretization schemes such as finite-element methods [7], finite volume methods [8], finite difference methods [9], among others. A critical computational consideration in simulating the impact and collisions of droplets is the treatment of the moving interface. This involves the solution of the adequate boundary conditions, coupled with the accurate representation and dynamic updating of the interface throughout the transient problem. Traditional mesh-based frameworks, commonly employed in the simulation of free surface or multiphase fluid flows, include the well-established front-tracking and front-capturing methodologies [10]. Beyond mesh-based approaches, mesh-free schemes have gained widespread application in simulating droplet dynamics. Smoothed Particle Hydrodynamics (SPH), for instance, offers an efficient and effective alternative of solving fluid dynamics problems without relying on a fixed grid allowing the solution of complex droplet interactions and deformations (see [11]). It is crucial to highlight the significance of both analytical papers within the context of impacting droplets [12] and experimental studies focusing on droplet dynamics [13]. These results play an important role in not only validating but also enriching the fidelity of computational simulations. Analytical works provide essential theoretical foundations, offering insightful equations and principles that guide numerical model development, as for instance the spreading of the droplet in a substrate [14]. Simultaneously, experimental investigations have been adopted to establish different scaling laws in the spreading of droplets [15, 16, 17, 18, 4, 19] and in droplet collisions [20, 21, 22, 23, 24]. The synergy between these analytical, experimental, and computational methods enhanced the understanding of droplet interactions [25, 26, 27].
More recently, the employment of Machine Learning (ML) strategies in the context of droplet dynamics received more attention due to the advances of new computation technologies, the increasing data availability and the implementation facilities on open-source codes (e.g. TensorFlow, Pytorch, MXNet, etc). The ML-based methods open a promising avenue to reduce both time and costs associated with traditional numerical simulations. This approach not only introduces efficiency gains but also serves as a robust alternative for predicting outcomes, reflecting the growing synergy between data-driven machine learning and traditional computational fluid dynamics. For example, in the context of droplet dynamics, different machine learning models have been adopted in order to predict physical or geometrical properties observed in complex flows, such as predicting the maximum spreading of a droplet by using the dimensionless numbers and properties related to each regime [28, 29, 30, 31, 32, 33]. The idea of applying related features to predict fluid properties and/or characteristics was also explored by predicting drop morphologies adopting image data [34], classifying splashing and nonsplashing drops using images from when half of the drop impacted the surface [35], using an explainable artificial intelligence (XAI) to interpret how said classifier works and uncover a possible correlation between the spreading of a drop and the normalized impact force on a solid surface [36], predicting interface curvature through neighboring particles from 2D front-tracking representation [37], predicting size, velocity and shape characteristics of droplets from droplet-on-demand liquid metal jetting (DoD-LMJ) [38], etc.
The connection between droplet dynamics and the energy budget has been extensively documented. When a droplet approaches a surface or interacts with another droplet, it initiates a process of energy transfer involving kinetic, surface, and dissipative components. Examining the energy budget in droplet dynamics not only enhances our understanding of fundamental fluid dynamics principles but also facilitates the development of simple analytical models based on energy balance. For instance, expressions derived from energy conservation can be employed to construct equations predicting the maximum diameter of a droplet spreading on a rigid surface. Seminal works by [15], [39], [40] and [19] exemplify the application of such energy balance principles.
Exploring the energy budget in droplet dynamics is of practical significance, offering valuable insights for optimizing processes such as inkjet printing. Achieving precise control over droplet impact dynamics is crucial for obtaining desired outcomes in applications like these. Recent literature has presented several papers delving into the energy budget in complex droplet dynamics. For instance, in the context of impinging droplets on rigid surfaces, [41] utilized a Volume-of-Fluid code to investigate the effects of low and high velocities on surface energy and viscous dissipation. Simultaneously, [42] conducted an energy-budget analysis using the diffusive-interface method, with a focus on a two-dimensional case. Additional insights into the connection between splash formation and energy dissipation were provided by [43], who employed experimental results involving a ferrofluid droplet with various shapes. In droplet collision studies, [44] applied a front-tracking numerical scheme to study the energy budget using unequal-size droplets while [45] combined an energy-based strategy and binary droplet collisions through experimental setups to reanalyze the bouncing regime boundary on the regime map proposed by [22]. More recently, [46] explored the energy balance during coalescence and breakup after collision, utilizing an experimental process for droplet atomization. Additionally, [47] investigated electro-mechanical energy conversion in droplets.
Therefore, based on the literature review described in the previous statement, it is evident that numerous studies have focused on various aspects, such as the maximum spreading diameter, different scaling laws, and boundaries on the regime maps of droplet collisions, using arguments of energy conservation. However, these results are typically presented in the context of static analyses, as there are few works [48, 41] that deal with the transient variation of properties (such as the spreading of a droplet impacting a surface or the diameter and height of a coalesced drop after a drop-drop collision) and their relation to the transient energy budget. Notably, recent investigations using learning algorithms are also limited to static data analysis. In this work, we propose a machine learning model based on a recurrent neural network to predict the temporal variation of the energy budget. We have adopted the Long Short-Term Memory (LSTM) [49] with a geometric input data set ${z(t_{1}),\dots,z(t_{n})}$, which learns a mapping to produce temporal predictions in terms of the energy budget. The term “geometric” refers to the time variation of the droplet diameter during spreading on a surface or the diameter and height of the coalesced droplet after the collision of two droplets. Expanding the potential of our framework utilizing LSTM, incorporating fluid parameters such as the Reynolds and Weber numbers, as well as flow parameters like impact parameter on the collision, enhances the accuracy of predictions for the transient behavior of kinetic, surface energies, and viscous dissipation. Although our tests have exclusively employed numerical results, the versatility of this non-intrusive methodology in predicting energy budgets during intricate droplet dynamics can be further extended to experimental investigations. This extension is feasible since it relies solely on temporal series data, specifically variations in droplet diameter, which can be easily captured by experimental devices.
@@ A.   Mathematical model and overview of the numerical method
The equations that govern the flow behavior of isothermal and incompressible fluids are the continuity (1) and momentum (2) equations:
@@
where $\mathbf{u}$ and $p$ are the velocity and pressure fields and $\rho$ is the fluid’s density. The stress tensor $\bm{\tau}$ assumes a form of a Newtonian liquid as
@@
using $U$ and $D$ as characteristic velocity and length, respectively. To simplify the nondimensionalization of the governing equations for isothermal and incompressible flows, we redefined them by using the previous dimensionless terms without its bars, as follows
@@
where $Re=\frac{\rho UD}{\eta}$ is the dimensionless Reynolds number.
In our tests, there are two main boundary conditions to be considered: solid walls condition, to define behavior across domain boundaries, and the free surface condition, for the contact between fluid and the external environment (air). In summary, a no-slip condition is applied for solid wall contours, while the free surface condition is defined as
@@
where $\bar{\kappa}=D\kappa$, $\mathbf{\bar{T}}=2\boldsymbol{\bar{\overset{.}{\gamma}}}-Re\cdot\bar{p}\mathbf{I}$ and $We=\frac{\rho U^{2}D}{\gamma}$ is the Weber number. Furthermore, an extra dimensionless number is adopted for the collision regime, the impact parameter $B=\frac{2b}{D_{1}+D_{2}}$, where $b$ is the center-to-center relative distance between the drops and $D_{1}$, $D_{2}$ are the diameters for each drop, as seen in [45].
To solve the governing equations numerically, we have employed the established projection method. In summary, the method uses the Helmholtz-Hodge decomposition to decouple velocity and pressure from equations (5) and (6). In the Helmholtz-Hodge decomposition, as presented in [50], any vector field defined in a region $\Omega$ with smooth boundary $\partial\Omega$ can be uniquely decomposed as
@@
with appropriated boundary conditions. After solving equation (14) and obtaining the velocity field $\tilde{\mathbf{u}}$, we still need to make sure the continuity equation will be satisfied. By using the Helmholtz-Hodge decomposition and equations (13) and (5), we can write the following Poisson equation for a scalar field $\psi$
@@
Finally, by solving equations (14) and (15) we obtain $\tilde{\mathbf{u}}$ and $\psi$ so that the decomposition (13) can be used to obtain a final velocity field $\mathbf{u}$ that satisfies both the momentum and continuity equations. The discretization process is achieved using the finite-difference method for free surface problems as well discussed in [51]. It is important to mention that we have employed an axisymmetric formulation to solve the droplet impacting problem, while droplet collision simulations were performed using a two-dimensional code.
@@ B.   Energy budget
Since there will be no gravitational influence on the generated data, the energy budget is given by the kinetic energy, the surface energy and the viscous dissipation. The kinetic energy $E_{k}$ is defined as
@@
calculated by the integral over the fluid volume V. Numerically, it is calculated as the sum of the kinetic energy inside all cells fully inside a droplet. The surface energy $E_{s}$ is computed as
@@
The law of conservation of energy defines that the total sum of all the energies ($E_{t}$) in the system must remain constant through time so that
@@
In order to construct a nondimesional investigation, we have adopted the following nondimensional energy quantities:
@@
The focus of this work is to use geometric time series data extracted from numerical simulations to generate time series predictions for each type of energy. From now on, we have omitted the bar over the nondimensional variables.
@@ C.   Problem descriptions and data preparation
To apply the neural network solution, we first need to generate adequate datasets. In order to do so, numerical simulations are used to generate two different flow regimes: drop impact on a solid surface, to generate diverse spreading lengths with different initial drop shapes, and droplet collision, to test the capacity of the model to predict features that oscillate through time. Both problem descriptions are represented in Fig. 1 (left) while the schematic to obtain the dataset is also described in the same figure (right).
@@
In order to generate the dataset, each case is executed with varying fluid and flow parameters, and/or initial drop shapes. In the droplet impact scenario, we vary the droplet shape by extracting interfaces from experimental droplet images [43]. Conversely, for the droplet collision problem, we adopt the impact parameter B to induce variations. In both cases, modifications are also conducted changing the fluid parameters based on Reynolds and Weber numbers. In summary, samples within the dataset can be defined as individual data points, each characterized by a specific combination of features encompassing fluid and flow parameters as well as initial drop characteristics, as it will be detailed in the next sections.
C.1 Spreading dataset
For the first dataset, we were inspired by the drop shapes from experiments presented by [43]. To test the efficiency of the machine learning model, the shapes were chosen in such a way that the train, validation and test subsets would have sufficiently different shapes, as shown in Fig. 2.
Fig. 2: Different initial drop shapes and the train-validation-test splits.
To extract particles from these shapes and simulate each case, an image interface extraction method was implemented, as described in Appendix A. The initial diameter and height of each shape, as presented in Table 1, are selected in such a way that each droplet has approximately the same volume. Since we used an axisymmetric simulation for the spreading case, as shown in Section A, this is necessary to make each droplet have approximately the same mass.
$Shape$
$D(0)$ (in mm)
$H(0)$ (in mm)
1
3.92
3.98
2
3.6771
4.64
3
2.62
4.64
4
4.64
3.6829
5
4.6086
3.1343
6
4.8143
2.5714
7
3.8229
4.24
8
4.5057
3.3857
9
3.5571
5.0086
10
4.6343
3.6229
Table 1: The initial spreading length and height for each drop shape.
The parameters chosen to configure the numerical simulations are presented in (Table 2). For each set of parameters and one droplet shape, a numerical simulation is run to generate one sample. Some combinations have fewer samples due to potential numerical errors in the solver, resulting in the failure to generate them. The Reynolds and Weber numbers were arbitrarily defined by varying the impact velocity. In total, 132 samples were extracted.

### v2

@@
Diego A. de Aguiar 111diego.aguiar@unesp.br, ORCID: $0$-0001-5159-867X Departamento de Matemática e Computação, Faculdade de Ciências e Tecnologia, Universidade Estadual Paulista “Júlio de Mesquita Filho”, Presidente Prudente, Brazil Hugo L. França 222franca.hugo1@gmail.com, ORCID: $0$-0002-5361-7704 Van der Waals-Zeeman Institute, Institute of Physics, University of Amsterdam, Amsterdam, The Netherlands Cassio M. Oishi333cassio.oishi@unesp.br, ORCID: $0$-0002-0904-6561 Departamento de Matemática e Computação, Faculdade de Ciências e Tecnologia, Universidade Estadual Paulista “Júlio de Mesquita Filho”, Presidente Prudente, Brazil
@@
The application of neural network-based modeling presents an efficient approach for exploring complex fluid dynamics, including droplet flow. In this study, we employ Long Short-Term Memory (LSTM) neural networks to predict energy budgets in droplet dynamics under surface tension effects. Two scenarios are explored: droplets of various initial shapes impacting on a solid surface and collision of droplets. Using dimensionless numbers and droplet diameter time series data from numerical simulations, LSTM accurately predicts kinetic, dissipative, and surface energy trends at various Reynolds and Weber numbers. Numerical simulations are conducted through an in-house front-tracking code integrated with a finite-difference framework, enhanced by a particle extraction technique for interface acquisition from experimental images. Moreover, a two-stage sequential neural network is introduced to predict energy metrics and subsequently estimate static parameters such as Reynolds and Weber numbers. Although validated primarily on simulation data, the methodology demonstrates potential for extension to experimental datasets. This approach offers valuable insights for applications such as inkjet printing, combustion engines, and other systems where energy budgets and dissipation rates are important. The study also highlights the importance of machine learning strategies for advancing the analysis of droplet dynamics in combination with numerical and/or experimental data.
keywords: Droplets |||| Numerical solution |||| Prediction |||| Surface Tension |||| LSTM |||| Energy budget
@@ I.   Introduction
Droplet dynamics is a complex phenomenon that covers multiple scientific and engineering domains, requiring a collaborative integration of experimental, analytical, and numerical approaches. Recent advances in these methodologies have increased our understanding of the impact of liquid droplets on rigid surfaces, allowing innovations in fields such as aerospace, energy, materials science, biotechnology, forensics, and the food and cosmetics industries. Notable contributions in this context can be found in [Alexander2006, Howard1994, Nikolopoulos2009, Christophe2016, Shiji2018]. These articles clarify some important aspects of this process such as: the dimensionless numbers and their influence on the maximum spreading of the droplet, different deformation regimes and their consequences in the dynamical energy budget of the system.
Analyzing the energy budgets in droplet dynamics is crucial for understanding fluid behavior and optimizing practical applications. When a droplet approaches a surface or interacts with another droplet, energy transfer occurs between the kinetic, surface, and dissipated components. Studying this interplay not only improves the understanding of fundamental fluid dynamics but also enables the improvement of models based on energy balance, such as those predicting the maximum spreading diameter of a droplet on a rigid surface [Chandra1991, Pasandideh1996, Nicolas1999, Chijioke2005, Vatsal2024]. Recent studies have advanced the understanding of energy-budget dynamics in droplet interactions, as for instance [Jae2016_2, Lennon2023, Qingzhe2021, Zhenyu2017, Karrar2019, Antonov2020, Antoine2021]. These works have practical relevance, particularly in applications as the inkjet printing process [Detlef2022], where precise control over droplet impact dynamics is essential for achieving desired outcomes. More specifically, the understanding of the collision process in inkjet printing is important, influencing the spatial distribution, size, and morphology of the printed droplets. As droplets traverse through air, dynamical fluid interactions can be observed, such as coalescence, fragmentation, bouncing and other complex behaviors. When droplets impact a substrate, factors such as velocity, impact angle, and material properties of both droplet and substrate influence the outcome. The subsequent spreading of droplets further complicates the process, governing the interaction between ink and substrate and affecting droplet shape [Jens2010]. These challenges highlight the intricate interplay of fluid dynamics in inkjet printing, including droplet collisions, substrate interactions, and spreading dynamics. The current work focuses on two key stages outlined in [Detlef2022]: droplet impact with varying shapes on a substrate and droplet coalescence during collisions.
Over the past decades, the simulation of droplet dynamics has been adopted for advancing our understanding of this complex fluid phenomena [Lorstad2004, Hua2015, Zhuan2022]. Computational frameworks employing the classical finite-element, finite-volume, or finite-difference discretization methods [Xing2007, Robert2010, Stephane2018] are combined with methods to track the deformable moving interface (e.g. front-tracking [Tryggvason2011]) and transient boundary conditions during droplet impact [Yu2005]. Simultaneously, experimental investigations have been adopted to establish different scaling laws in the spreading of droplets [Chandra1991, Nick2014, Fukai1995, Jae2016, Shiji2018, Chijioke2005] and in droplet collisions [Kuo2008, Estrade1999, Qian1997, Qian1997_2, Willis2000]. The synergy between these analytical, experimental, and computational methods enhanced the understanding of droplet interactions [Sander2016, Sikalo2005, Attane2007, Ilia2009].
More recently, the employment of machine learning-based methods [Song2022, Moussa2022, Jiguo2023, Lap2023, Elham2020, Raihan2023] for droplet dynamics opens a promising avenue to reduce both time and costs associated with traditional numerical simulations. Notably, most of these works are limited to static data analysis [Shihao2024, Moussa2022, Shengzhi2024, Jiguo2023, Lap2023, Liu2021, Hugo2022, Elham2020, Wenlong2024], with few exceptions as [Olinka2020, Jae2016_2]. Therefore, the motivation of this study is two-fold. Firstly, we propose a machine learning model based on Long Short-Term Memory (LSTM) [Hochreiter1997] to predict the temporal variation of the energy budget by incorporating dimensionless parameters. Second, through using a method designed for extracting particles from interfaces in digital images we extend our framework to be employed for experimental investigations. With advancements in open-source computational tools (e.g., OpenFOAM, Basilisk) and high-speed imaging techniques, our study highlights the potential of physics-informed machine learning models for predicting and analyzing droplet dynamics during transient regimes. While our current tests focus on numerical results for Newtonian fluids, the non-intrusive concept of this approach, adopting only temporal data, allows its applicability to non-Newtonian fluids. The accurate prediction of energy budgets in complex droplet dynamics demonstrates the robustness and versatility of our methodology, positioning it as a valuable tool for future research.
@@ A.   Mathematical model and overview of the numerical method
The equations that govern the flow behavior of isothermal and incompressible fluids are the continuity and momentum equations, respectively:
@@
where $\mathbf{u}$ and $p$ are the velocity and pressure fields and $\rho$ is the fluid density. The stress tensor $\bm{\tau}$ assumes the form of a Newtonian liquid as
@@
using $U$ and $L$ as characteristic velocity and length, respectively. Dropping the bars, for convenience, the dimensionless version of equations (1)-(2) is
@@
where $Re=\frac{\rho UL}{\eta}$ is the dimensionless Reynolds number.
In our tests, two main boundary conditions are considered: no-slip boundaries, to simulate the impact of a droplet onto a rigid plate, and the free surface condition, for the interface between fluid and the external environment (air). This free surface condition is defined as
@@
where $\bar{\kappa}=L\kappa$, $\mathbf{\bar{T}}=2\boldsymbol{\bar{\overset{.}{\gamma}}}-Re\cdot\bar{p}\mathbf{I}$ and $We=\frac{\rho U^{2}L}{\gamma}$ is the Weber number. In the case of collision between two droplets, an additional dimensionless number is adopted, the impact parameter $B=\frac{2b}{D_{1}+D_{2}}$, where $b$ is the center-to-center relative distance between the droplets and $D_{1}$, $D_{2}$ are the diameters for each droplet, as seen in [Karrar2019].
To solve the governing equations numerically, we have employed the established projection method, utilizing the Helmholtz-Hodge decomposition to decouple velocity and pressure from equations (5)-(6). In this decomposition, presented in [Chorin1979], any vector field defined in a region $\Omega$ with smooth boundary $\partial\Omega$ can be decomposed as
@@
with appropriate boundary conditions. After solving equation (14) and obtaining the velocity field $\tilde{\mathbf{u}}$, it is still necessary to make sure the continuity equation is satisfied. By combining equations (13) and (5), the following Poisson equation can be written
@@
Finally, by solving equations (14) and (15) we obtain $\tilde{\mathbf{u}}$ and $\psi$ so that the decomposition (13) can be used to obtain a final velocity field $\mathbf{u}$ that satisfies both the momentum and continuity equations. Equation (15) is solved with homogeneous Neumann boundary condition on rigid walls, while a third type of boundary condition is used for $\psi$ at the free surface to improve numerical stability. More details about the mathematical formulation and boundary conditions in this method can be found in [Oishi2019]. The discretization process is achieved using the finite-difference method for free surface problems, as detailed in [Oishi2016]. An axisymmetric formulation was adopted for the droplet impact simulations, while droplet collision simulations were performed using a two-dimensional code.
@@ B.   Energy budget
Without gravity, the total energy budget is given by the kinetic energy, surface energy and viscous dissipation. The kinetic energy $E_{k}$ is defined as
@@
obtained by integrating over the fluid volume $V$. Numerically, it is calculated as the sum of the kinetic energy inside all cells fully inside a droplet. The surface energy $E_{s}$ is computed as
@@
Conservation of energy states that the total sum of all the energies ($E_{t}$) in the system must remain constant through time, that is
@@
In order to keep our investigation nondimensional, we also rescale the energies (16)-(18) and work with the following quantities
@@
The focus of this work is to use geometric time series data extracted from numerical simulations to generate time series predictions for each energy component. From now on, we omit the bars over all nondimensional variables for convenience.
@@ C.   Problem descriptions and data preparation
To apply a neural network, it is essential to create a sufficiently representative training dataset. For this purpose, numerical simulations are employed to generate two different flow regimes: (i) droplet impact on a solid surface, to generate diverse spreading diameters with different initial droplet shapes, and (ii) binary droplet collision, to test the capacity of the model to predict features that oscillate through time. Both problem descriptions are represented in Fig. 1 (left) while the schematic to obtain the dataset is also described in the same figure (right). Further details about the codes used to generate the datasets can be found in [Oishi2019, Hugo2022_2].
@@
In order to generate the datasets, each case is executed with varying flow parameters, and/or initial droplet shapes. In the droplet impact scenario, we vary the droplet shape by extracting interfaces from experimental images [Qingzhe2021]. For binary droplet collisions, we induce initial geometrical variations through the impact parameter B. In both cases, we also vary the Reynolds and Weber numbers. In summary, samples within the dataset can be defined as individual data points, each characterized by a specific combination of features encompassing dimensionless parameters ($Re$ and $We$) with initial geometrical droplet characteristics, as will be detailed in the following sections.
C.1 Droplet spreading dataset
Our first dataset is inspired by droplet shapes from experiments presented by [Qingzhe2021]. To test the efficiency of the machine learning model, the shapes were selected to ensure that the training, validation and test subsets would have sufficiently different shapes, as shown in Fig. 2.
Fig. 2: Different initial droplet shapes and the train-validation-test splits. The more complex shapes were specifically chosen for the validation and test sets.
An interface extraction method was implemented in order to obtain these shapes from experimental images, as described in Appendix A. The initial diameter and height of each experimental shape, as presented in Table 1, are selected in such a way that each droplet has approximately the same volume.
Shape $D(0)$ (in mm) $H(0)$ (in mm)
1 3.92 3.98
2 3.6771 4.64
3 2.62 4.64
4 4.64 3.6829
5 4.6086 3.1343
6 4.8143 2.5714
7 3.8229 4.24
8 4.5057 3.3857
9 3.5571 5.0086
10 4.6343 3.6229
Table 1: The initial spreading diameter and height for each droplet shape.
The parameters chosen to configure the numerical simulations are presented in Table 2. For each set of parameters and one droplet shape, a numerical simulation is run to generate one sample. The Reynolds and Weber numbers were arbitrarily defined by varying the impact velocity, with the characteristic length $L$ defined as the initial droplet diameter $D(0)$. In total, 132 simulations were performed, each run until a maximum time of $t_{\text{max}}=50$, with 1000 equally spaced snapshots extracted per simulation.